# Benchmarking Fake Voice Detection in the Fake Voice Generation Arms Race

### v1

@@ 4.3 Cross-Domain Evaluation Performance Analysis
We perform a comprehensive analysis based on Table Table 6 and 7 with three focus. In Section 4.3.2, we discuss one generator performance across detector. In Section 4.3.3, we discuss one detector performance across different generator. Then finally, we provide our overall score analysis in Section 4.3.4 from the both ways to evaluate the overall detector performance.
Table 6: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 1. Column best (lowest) values are blue, worst (highest) red. All metrics are the lower the better.
@@
4.3.1 Performance Metric for Fake Voice Detectors
Following the ASVSpoof Challenge methodology (e.g., [125]), we adopted the minimum normalized Detection Cost Function (minDCF) and Equal Error Rate (EER) as primary metrics for evaluating countermeasure systems. Additionally, the Log Likelihood Ratio Cost ($C_{llr}$) is often used to evaluate not only discrimination but also the calibration of detector scores. To emphasize the overall fake voice detector performance across various generators while also considering model efficiency, we propose a novel Detector Overall Score ($S_{ovrl}$).
We computed a single unified overall score ($S_{i}$) for each fake voice detector $i$. This score combines the detector’s empirical performance against multiple fake voice generators with a penalty for its model size. The construction is as follows:
First, raw performance metrics for detector $i$ against generator jj—EER ($m_{ij}^{(\text{EER})}$), $C_{llr}$ ($m_{ij}^{(C_{\text{llr}})}$), and minDCF ($m_{ij}^{(\text{minDCF})}$), generically denoted $m_{ij}$—were linearly rescaled to $[0,1]$. This normalization uses the global minimum ($m_{\text{min}}$) and maximum ($m_{\text{max}}$) values for that specific metric type:
$n_{ij}=\begin{cases}\frac{m_{ij}-m_{\text{min}}}{m_{\text{max}}-m_{\text{min}}},&\text{if }m_{\text{max}}\neq m_{\text{min}},\\ 0,&\text{otherwise.}\end{cases}$
This yields normalized metrics $n_{ij}^{(\text{EER})}$, $n_{ij}^{(C_{\text{llr}})}$, and $n_{ij}^{(\text{minDCF})}$, where lower values indicate better performance.
Next, to emphasize challenging fake voice generation methods (e.g., Text-to-Speech (TTS), TTS with Voice Conversion, Audio Reconstruction), their contributions were weighted. For each generator $j$, a challenge score $R_{j}$ was assigned, with higher $R_{j}$ for more challenging generators. These $R_{j}$ scores were normalized to produce weights $w_{j}$:
$w_{j}=\frac{R_{j}}{\sum_{k}R_{k}}$
where the sum is over all generator challenge scores. For example, if for three generators $G_{1},G_{2},G_{3}$, their assigned scores are $R_{1}=2,R_{2}=1,R_{3}=1$ respectively (total $\sum_{k}R_{k}=4$), the weights become $w_{1}=0.5,w_{2}=0.25,w_{3}=0.25$. This ensures more difficult generators receive greater impact.
The average score $A_{ij}$, detector $i$’s combined performance against generator $j$, is the mean of its three normalized metrics:
$A_{ij}=\frac{1}{3}\left(n_{ij}^{(\text{EER})}+n_{ij}^{(C_{\text{llr}})}+n_{ij}^{(\text{minDCF})}\right)$
Overall empirical performance $P_{i}$ for detector $i$ is the weighted sum of $A_{ij}$ scores over all generators $j$:
$P_{i}=\sum_{j}w_{j}A_{ij}$
For model complexity, parameter counts $C_{i}$ were log-transformed ($\log C_{i}$) and normalized. Let $L_{\text{min}}=\min_{k}\log C_{k}$ and $L_{\text{max}}=\max_{k}\log C_{k}$. The normalized log-parameter count $\hat{C}_{i}$ is:
$\hat{C}_{i}=\begin{cases}\frac{\log C_{i}-L_{\text{min}}}{L_{\text{max}}-L_{\text{min}}},&\text{if }L_{\text{max}}\neq L_{\text{min}},\\ 0,&\text{otherwise.}\end{cases}$
This maps $\log C_{i}$ to $\hat{C}_{i}\in[0,1]$, where higher values mean larger, less efficient models.
Finally, the overall score $S_{i}$ blends empirical performance $P_{i}$ and model complexity penalty $\hat{C}_{i}$ using a trade-off factor $\alpha\in[0,1]$:
$S_{i}=\alpha P_{i}+(1-\alpha)\hat{C}_{i}$
The resulting $S_{i}$ is a consolidated, lower-is-better score, reflecting both detector accuracy and efficiency.
@@ 4.3.2 Generalization Analysis of Fake Voice Generators
This section analyzes how state-of-the-art detectors generalize across different families of fake voice generators. We discuss performance trends using the Equal Error Rate (EER, %) metrics reported in Table 6 and 7.
Aggregated Datasets. While detectors perform well on standard benchmarks like ASVSpoof-21LA, with top systems achieving EERs below $2\%$ (e.g., $0.82\%$ for XLSR-AASIST), their performance degrades significantly under distribution shifts. EERs increase on compressed audio like ASVSpoof-21DF (e.g., XLSR-AASIST to $2.85\%$) and data with waveform degradations in FoR (XLSR-SLS at $4.07\%$). The challenge is most severe in cross-lingual settings, such as the Chinese CFAD dataset, where even robust models like XLSR-SLS see EERs rise to $13.06\%$, and weaker ones like AASIST approach chance-level performance ($49.11\%$). This highlights a critical weakness in out-of-distribution generalization.
@@
Table 7: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 2. Column best (lowest) values are blue, worst (highest) red. All metrics are the lower the better.
@@
Neural Codec Language Model TTS This family, including MaskGCT-TTS and FireRedTTS-1S, represents the most difficult detection scenario. These models leverage a neural audio codec as a final polishing stage, explicitly optimizing for perceptual realism. Their advanced techniques—such as mask-predict paradigms, two-step decoding, and zero-shot voice conversion—produce speech with fewer artifacts and more natural prosody. Consequently, detector performance is severely compromised. The best EER against MaskGCT-TTS is $24.94\%$ (XLSR-SLS), and it rises to $30.86\%$ for the even more advanced FireRedTTS-1S (XLSR-Conf.+TCM). For these models, many detectors fail entirely, with EERs exceeding $60\%$.
@@ 4.3.3 Generalization Analysis of Fake Voice Detectors
End-to-End Models End-to-end models generally exhibit limited generalization, characterized by high error rates and poor calibration when faced with diverse and unseen speech generators. Performance within this category is inconsistent; for instance, RawNet2 often struggles across most generators, while RawPCDart can excel against specific types. A common weakness is unstable calibration, with models like RawNet2 and RawBMamba exhibiting high $C_{llr}$ scores.
This brittleness arises because standard end-to-end training objectives (e.g., cross-entropy) do not explicitly enforce robustness to domain shifts. The convolutional front-ends used in models like RawNet2 [113] and RawBMamba [18] learn features that are highly discriminative for in-domain data but fail to generalize to novel synthesis artifacts. Even AASIST [45], despite its improved graph attention classifier, is ultimately constrained by its RawNet2-based encoder and shows more generalization failures than detectors with more specialized architectures.
Wav2Vec-2.0 Feature Extratcion Models This category of models demonstrates highly context-dependent performance, achieving state-of-the-art results against certain generator families but failing significantly on others. For example, XLSR-AASIST excels on in-domain data like ASVSpoof-21LA ($0.82\%$ EER) but is less effective against modern TTS systems. This sensitivity to generator type is a defining characteristic: performance often collapses against advanced unseen systems (e.g., Matcha-TTS, FireRedTTS-1S) while remaining strong against older GAN or diffusion models. Among these models, XLSR-Conformer+TCM often demonstrates superior score calibration.
@@ 4.3.4 Detector Overall Performance
Conversely, very lightweight models like AASIST ($S_{i}=0.5053$) and RawBMamba ($S_{i}=0.5681$) attained competitive mid-tier scores. Their minimal $\hat{C}_{i}$ penalties compensated for what Section 4.3.3 described as more modest or variable empirical performance (e.g., "limited generalization" or higher $C_{llr}$ for some end-to-end models). Detectors such as RawPCDart ($S_{i}=0.7739$) and RawNet2 ($S_{i}=0.7030$) ranked lower, as their less competitive $P_{i}$ scores were not sufficiently counterbalanced by their model complexities. These $S_{i}$ rankings thus provide a holistic but weight-dependent view, confirming that while the newest generator architectures pose significant detection hurdles (in Section 4.3.2), detectors excelling in broad empirical accuracy can lead when model efficiency is a secondary but still influential consideration.
@@ 5 Future Directions
In this section, we outline concrete and forward-looking research avenues to navigate the escalating arms race between synthetic voice generation and detection, highlighting needs of both technical innovations and systemic strategies.
@@ 6 Conclusion
Appendix A Ethical Considerations
The development and evaluation of fake voice generation and detection technologies raise critical ethical concerns. While progress in synthetic voice technologies supports beneficial applications such as accessibility for speech-impaired individuals, educational tools, and creative content production, it simultaneously creates risks for fraud, impersonation, misinformation, and erosion of social trust. Our research acknowledges this dual-use nature and adopts a framework grounded in The Menlo Report principles—Beneficence, Respect for Persons, Justice, and Respect for Law and Public Interest.
Stakeholders.
Key stakeholders include (i) the general public, who may benefit from stronger detection systems but could be harmed by misuse of synthetic voices; (ii) companies and organizations that rely on secure authentication and communication systems; (iii) the research community, which benefits from reproducibility and benchmarking; and (iv) the research team itself, which must ensure its work does not contribute to misuse.
Harms and Risks.
Two broad categories of potential harms are considered: tangible harms and violations of rights. Tangible harms include financial loss, reputational damage, or security breaches stemming from voice spoofing. Rights-based harms include the violation of privacy expectations, unauthorized use of personal voice likeness, and erosion of informed consent in data collection. Additionally, synthetic voices could exacerbate social harms such as misinformation campaigns or targeted harassment.
Mitigations.
We proactively mitigate these risks by: (1) restricting experimentation to publicly available datasets and pre-trained models; (2) avoiding the release of maliciously usable or identity-targeted audio; (3) limiting all generated content strictly to controlled benchmarking and analysis; (4) not deploying generated audio in real-world systems; and (5) encouraging safeguards such as controlled access, watermarking, and forensic traceability in downstream use. Risks that cannot be fully eliminated, such as the potential misuse of broader methodological insights, are acknowledged and documented transparently.
Decision to Proceed.
We judged that the ethical benefits—improving the robustness of fake voice detection and highlighting system vulnerabilities—outweigh the risks. This aligns with the Beneficence principle by maximizing potential societal benefit while minimizing foreseeable harm. Our analysis also respects the rights of individuals by not using private or sensitive data and by ensuring research outcomes are oriented toward defensive, rather than offensive, applications.
Responsible Disclosure and Community Norms.
Where vulnerabilities in detection systems are identified, we follow responsible disclosure practices, providing advance notice to relevant stakeholders (e.g., detection model developers) before public release of findings. This approach reduces potential negative outcomes while ensuring that security improvements are actionable.
Researcher Wellbeing.
We also acknowledge that exposure to adversarial audio datasets can impose cognitive or psychological strain on research team members. Appropriate measures, such as workload distribution and technical filters, were implemented to mitigate these risks.
Conclusion.
Our work is conducted with a commitment to both ethical integrity and legal compliance. While no research of this nature is without risk, our structured approach to stakeholder analysis, harm mitigation, and responsible dissemination provides a defensible ethical justification for both the conduct of this research and its publication. We encourage future researchers to build upon our work with equal or greater attention to ethical responsibility.
Appendix B Open Science
In order to advance transparency, reproducibility, and community-driven research, we are committed to open science practices. Due to the large size of the benchmark dataset used in our cross-domain evaluation, we provide a partial release at the following link: https://1drv.ms/u/c/e21f0906cf1459ef/Ef8uGy3nQdNFhYBBGKBQlFQBiOCOfoYFrN3JK6-z0iGy6A?e=6dccQU. We also have evaluation code in https://anonymous.4open.science/r/benchmark_voice-256F/ for fake voice detection which forked from ASVSpoof-5 we used in cross evaluation (including our composite generation and detection metric evaluation). Upon publication, we will release the complete and fully curated dataset and code to the community to ensure unrestricted access and facilitate future research.
Appendix C Related Work
Generalization and Cross Evaluation. Performance often collapses when detectors face unseen generators or corpora [81, 105, 34, 32]. Doan et al. [24] identify highly transferable artifacts from large E2E-GAN generators, and Mishra et al. [79] use Shapley-based explanations to pinpoint shortcut features. Wu et al.’s CodecFake [135] confirms that retraining on neural codec data enhances cross-dataset resilience.
Appendix D Fake Voice Generator
Table 9: An overview of additional TTS models, supplementing Table 1. Modeling: Trans=Transformer, Diff=Diffusion, AR=Autoregressive, VAE=Variational Autoencoder, VQ=Vector Quantization, GAN=Generative Adversarial Network, FM=Flow Matching. Type: AM=Acoustic Model, Fully E2E=Fully End-to-End, NC-LM=Neural Codec Language Model. Params (M): Model parameters in millions. Open Src: ✓=Open Source, $\times$=Not Open Source. Control (Prosody/Style): ✓=Supported, $\times$=Not Supported. Dataset Spk: =Single Speaker, =Multi-Speaker (2-1000), =Large Multi-Speaker (1000+). Dataset Hours: =<50 Hours, =50-100 Hours, =>100 Hours. Dataset Languages: =Single Language, =Few Languages (2), =Multilingual (3+). Inference Speed: =Very fast (10x+ real-time), =Fast (2-10x), =Real-time or slower. Zero/One Shot: ✓=Supports zero-shot, $\times$=No voice conversion. Values are based on original papers. – indicates value not stated.
Paper Year Type Modeling Params Open Control Dataset Inference Zero
(M) Src Prosody Style Spk Hours Languages Speed Shot
Delightful-TTS 2021 AM Trans $\times$ $\times$ $\times$
HierSpeech++ 2022 NC-LM VAE 97
Matcha-TTS 2023 AM FM 18.2 $\times$ $\times$ $\times$
Mega-TTS 2023 Fully E2E VQ-GAN 222.5 $\times$
UniCATS 2023 NC-LM VQ-Diff+Trans $\times$
VoiceCraft 2024 NC-LM Trans 830
E2-TTS 2024 AM Trans + FM 335
CLaM-TTS 2024 NC-LM VAE $\times$
FireRedTTS 2024 NC-LM Trans 400
X-TTS 2024 NC-LM VQ-VAE+Trans 482
CosyVoice 2024 NC-LM FM+Trans
Table 9 provides an extended overview of recent TTS models beyond those in the main table, highlighting their modeling approaches (e.g., Transformer, Diffusion, VAE, GAN, Flow Matching), type (acoustic, end-to-end, or neural codec LM), parameter sizes, availability of open-source implementations, prosody/style control, dataset scale in terms of speakers, hours, and languages, as well as inference speed and zero/one-shot capability. This supplementary comparison illustrates the diversity of design choices in fake voice generation systems.
Table 10: Metrics, normalization range, and aggregation weights.
Task Metric Range Weight
TTS DNSMOS $(0,1]$ $\alpha_{1}=0.25$
SIM $(0,1]$ $\alpha_{2}=0.25$
PSNR $(0,1]$ $\alpha_{3}=0.15$
WER $(0,1]$ $\alpha_{4}=0.20$
RTF $(0,1]$ $\alpha_{5}=0.075$
LogParams $(0,1]$ $\alpha_{6}=0.075$
Audio Recon. DNSMOS $(0,1]$ $\beta_{1}=0.25$
PESQ $(0,1]$ $\beta_{2}=0.225$
PSNR $(0,1]$ $\beta_{3}=0.15$
SSIM $(0,1]$ $\beta_{4}=0.225$
RTF $(0,1]$ $\beta_{5}=0.075$
LogParams $(0,1]$ $\beta_{6}=0.075$
Appendix E Fake Voice Datasets
Fake voice datasets are the cornerstone of developing and evaluating fake voice detection systems, and the chosen algorithms used to generate the fake voices in the dataset define its quality. The composition and characteristics of the real and generated data profoundly influence detector design, performance, and ultimately, their generalization capabilities. Table 11 presents an overview of commonly used and recent novel datasets for training, development and evaluation. We also note the generator types associated with each fake voice dataset.
Table 11: Overview of fake voice detection datasets. Audio condition: Type of audio degradation present (clean, noisy or codec). #spoof:#bona: Approximate number of spoofed vs. bonafide utterances (k=thousands, M=millions). Unseen attack: ✓=Contains attacks/generators potentially unseen during system training phase, $\times$=All attacks/generators are known during training evaluation setup. Language: EN=English, ZH=Chinese, JP=Japanese, Multiple=more than one language. Generator type: Synthesis or transduction methods used for creating spoofing attacks (✓=Present). Sub-columns: DSP=Digital Signal Processing, AR=Autoregressive models, GAN=Generative Adversarial Network-based, Diff=Diffusion-based, Flow=Flow-based, NCodec=Neural Codec based, F-E2E=Fully End-to-End models, Trans=Transformer-based. Empty entry means not specified or not mentioned in the dataset paper.
Dataset Year Condition #spoof:#bona Unseen Language Generator type
attack DSP AR GAN Diff Flow NCodec F-E2E Trans
ASVspoof-19LA[129] 2019 Clean 90k:10k EN
FoR [94] 2019 Clean 87k:108k EN
ASVspoof-21LA[73] 2021 Codec 133k:15k EN
ASVspoof-21DF[73] 2021 Codec 520k:15k EN
Wavefake [29] 2021 Clean 118k:0 $\times$ EN, JP
Into The Wild [81] 2022 Noisy 12k:20k EN
CFAD [76] 2023 Noisy, Codec 77k:39k ZH
LibriSeVoc [110] 2023 Clean 79k:13k $\times$ EN
DFADD [25] 2024 Clean 164k:44k $\times$ EN
CVoiceFake [70] 2024 Codec 1.25M:0 $\times$ Multiple
ASVspoof-5 [125] 2024 Noisy, Codec 288k:1.21M EN
CodecFake [135] 2024 Codec 44k:44k $\times$ EN
CodecFake+ [17] 2025 Codec 90k:1.42M EN
Pivotal datasets in the field, particularly those from the ASVspoof challenge series (ASVspoof19, ASVSpoof-21, ASVspoof-5) [129, 73, 125], have been instrumental in standardizing evaluation protocols and driving progress. These challenges typically define specific tasks (e.g., Logical Access focusing on TTS and voice conversion and Physical Access on replay) and provide curated training, development, and evaluation sets. Other datasets like Fake or Real (FoR) [94] and the Chinese Fake Audio Detection (CFAD) [76] offer variations in languages, synthesis methods, or specific conditions. However, those datasets have some limitations: (1) Techniques Coverage Lag: The rapid evolution of voice generation techniques (the arms race described in Section 2) means datasets often lag behind, lacking examples from the latest generators. (2) Aggregation Masking: Traditional evaluations often aggregate diverse fake voices into a single evaluation set. This masks vulnerabilities to specific fake voices types.
Appendix F Fake Voice Evaluation Metrics
Audio was processed at 16kHz where specifically required in PESQ and DNSMOS.
  • TTS Evaluation:
    • Intelligibility: We fed each generated utterance into OpenAI’s Whisper ASR model (medium-en variant) to produce a hypothesis transcript $\hat{T}$ and compare it to the reference transcript TT.[90].
    • Speaker Similarity: For SIM, we used the pretrained WavLM Base+ model [14] to extract deep audio representations. This SIM value is compare with the ground truth same audio clip as a pairing with the fake voice.
    • Perceptual Quality: DNSMOS (overall score) by using official pip package.
    • Spectrum and Signal Fidelity: Peak Signal-to-Noise Ratio (PSNR) quantifies how closely the generated audio’s spectrum matches the reference. We computed the magnitude spectra $X_{\mathrm{ref}}[n]$ and $X_{\mathrm{test}}[n]$ via STFT and normalized both so that the maximum spectral magnitude is 1.
  • Audio Reconstruction Evaluation:
    • Perceptual Quality: PESQ(official), DNSMOS(same as TTS evaluation).
    • Spectrum and Signal Fidelity: SSIM (STFT-based), LS-MSE (log Mel-Spectrograms comparison): we extracted the 80-dimensional mel-spectrogram features using 40 ms Hanning window, 12.5 ms frame shift, 1024-point FFT, and 0 Hz & 12 kHz lower & upper frequency cutoffs following [2], PSNR (same as in TTS evaluation).
  • Computational Complexity:
    • Parameters: We checked the parameters of each models based on their provided checkpoints in their official repository. For Hifi-GAN, MelGAN and ParallelWaveGAN, we used the provided checkpoints from https://github.com/kan-bayashi/ParallelWaveGAN and For FastSpeech2 checkpoints with Hifi-GAN, we used the provided checkpoints in the Google Colab implementation in https://github.com/espnet/espnet.
    • Inference Speed: We use Real-time factor (RTF) to measure the inference speed which can be calculated by the ratio of the time it takes to generate a fake voice to the total duration of this audio, and then aggregated and divided all clips RTFs to directly calculate a mean average value. For text-to-speech generators (including TTS with voice conversion), we include the vocoding time in consideration.
Appendix G Evaluation Metric Normalization
All evaluation metrics are normalized to the open interval $(0,1]$ prior to aggregation in the overall score formulas (Eq. 1–2). Normalization ensures comparability across heterogeneous scales. Let $x$ denote the raw metric value, then the normalized value is given by:
$M_{\text{norm}}=\frac{x-x_{\min}}{x_{\max}-x_{\min}}\cdot(1-\epsilon)+\epsilon,\quad\epsilon>0,$ (1)
where $x_{\min}$ and $x_{\max}$ are the lower and upper bounds for the metric across the evaluation set, and $\epsilon$ is a small constant (e.g., $0.1$) ensuring $M_{\text{norm}}\in(0,1]$. The weights are shown in Table 10.
Appendix H Limitations
  • Model Coverage: Our benchmark evaluates 20+ generators and 8 detector architectures. Despite this breadth, the field moves rapidly—multiple neural codec TTS or generalizable fake voice detector released after 2025 are absent. Consequently, the reported rankings represent a time-stamped evaluation rather than an exhaustive leaderboard.
  • Dataset & Environment: All experiments use clean, single-speaker English downloaded online as original dataset. We do not test noisy or codec audio or low-resource languages whose phonotactics differ markedly. Generalization of both generators and detectors to such conditions therefore remains unverified.
  • ASV Consideration: We focus on fake voice detection countermeasure, but do not measure how each fake voices interferes with automatic speaker verification (ASV) systems. Generator-specific artifacts could degrade ASV scores depending on the back-end. A unified evaluation linking fake voice detectability with ASV system is needed in future work.

### v2

@@ 4.3 Cross-Domain Evaluation Performance Analysis
We perform a comprehensive analysis based on Table 5 and Table 6 with three foci. In Section 4.3.2, we discuss one generator performance across detector. In Section 4.3.3, we discuss one detector performance across different generator. Then finally, we provide our overall score analysis in Section 4.3.4 from the both ways to evaluate the overall detector performance.
Table 5: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 1. The best and worst performing values for each metric are highlighted in blue and red, respectively. All metrics are the lower the better.
@@
4.3.1 Performance Metrics for Fake Voice Detectors
We adopt standard metrics from the ASVSpoof Challenge: Equal Error Rate (EER) for discrimination, and minimum normalized Detection Cost Function (minDCF) and Log Likelihood Ratio Cost ($C_{llr}$) for a combination of discrimination and score calibration. To synthesize these results, we propose a Detector Overall Score ($S_{i}$) that balances empirical performance against model complexity.
The score for detector $i$ is calculated as $S_{i}=\alpha P_{i}+(1-\alpha)\hat{C}_{i}$, where $P_{i}$ is the normalized empirical performance and $\hat{C}_{i}$ is the normalized model complexity penalty (Parameter). The empirical performance $P_{i}$ is a weighted average of the detector’s performance against each generator $j$, calculated as $P_{i}=\sum_{j}w_{j}A_{ij}$, where $A_{ij}$ is the average normalized EER, minDCF, and $C_{llr}$.
Justification of Parameters: Here we provide a rationale for our parameter choices:
  • Generator-Specific Weights ($w_{j}$): The weights $w_{j}$ are designed to ensure a detector’s final score is more significantly impacted by its performance against more evasive, high-quality generators. To achieve this systematically, the weight for each generator $j$ is derived from its Challenge Score ($R_{j}$). We define this score based on a direct principle: the higher the quality of the synthetic voice, the more challenging it is to detect. Therefore, a generator’s Challenge Score ($R_{j}$) is set to be directly proportional to its overall quality score ($F_{\text{TTS}}$ or $F_{\text{Audio Reconstruction}}$) that we calculated in the previous section. This method ensures that generators producing the most realistic audio, such as the high-scoring neural codec LMs, are assigned the greatest weight, thereby focusing the evaluation on the most potent security threats.
  • Trade-off Factor ($\alpha$): We set $\alpha=0.8$ to heavily prioritize a detector’s empirical performance ($P_{i}$) over its model size penalty ($\hat{C}_{i}$). In the security context of fake voice detection, effectiveness is paramount, while model efficiency is a secondary, albeit important, consideration for practical deployment. This choice reflects that priority.
This scoring methodology provides a transparent, albeit opinionated, framework for ranking detectors based on a holistic view of their robustness and efficiency.
@@ 4.3.2 Generalization Analysis of Fake Voice Generators
This section analyzes how State-Of-The-Art detectors generalize across different families of fake voice generators. We discuss performance trends using the Equal Error Rate (EER, %) metrics reported in Table 5 and 6.
Aggregated Datasets. While detectors perform well on standard benchmarks like ASVSpoof-21LA, with top systems achieving EERs below $2\%$ (e.g., $0.82\%$ for XLSR-AASIST), their performance degrades significantly under distribution shifts. EERs increase on compressed voice like ASVSpoof-21DF (e.g., XLSR-AASIST to $2.85\%$) and data with waveform degradations in FoR (XLSR-SLS at $4.07\%$). The challenge is most severe in cross-lingual settings, such as the Chinese CFAD dataset, where even robust models like XLSR-SLS see EERs rise to $13.06\%$, and weaker ones like AASIST approach chance-level performance ($49.11\%$). This highlights a critical weakness in out-of-distribution generalization.
@@
Table 6: Performance Individual metrics (EER(%), $C_{llr}$, minDCF) - Part 2. The best and worst performing values for each metric are highlighted in blue and red, respectively. All metrics are the lower the better.
@@
Flow Matching TTS. Flow matching systems like Matcha-TTS pose a formidable challenge by modeling real data distributions with high fidelity. This technique effectively evades most detection mechanisms, resulting in poor detector performance. The lowest EER achieved is $24.29\%$ (RawBMamba), while many sophisticated detectors perform near chance level, with EERs approaching $50\%$.
Neural Codec Language Model TTS This family, including MaskGCT-TTS and FireRedTTS-1S, represents the most difficult detection scenario. These models leverage a neural audio codec as a final polishing stage, explicitly optimizing for perceptual realism. Their advanced techniques—such as mask-predict paradigms, two-step decoding, and zero-shot voice conversion—produce voice with fewer artifacts and more natural prosody. Consequently, detector performance is severely compromised. The best EER against MaskGCT-TTS is $24.94\%$ (XLSR-SLS), and it rises to $30.86\%$ for the even more advanced FireRedTTS-1S (XLSR-Conf.+TCM). For these models, many detectors fail entirely, with EERs exceeding $60\%$.
@@ 4.3.3 Generalization Analysis of Fake Voice Detectors
End-to-End Models End-to-end models generally exhibit limited generalization, characterized by high error rates and poor calibration when faced with diverse and unseen fake voice generators. Performance within this category is inconsistent; for instance, RawNet2 often struggles across most generators, while RawPCDart can excel against specific types. A common weakness is unstable calibration, with models like RawNet2 and RawBMamba exhibiting high $C_{llr}$ scores.
This brittleness arises because standard end-to-end training objectives (e.g., cross-entropy) do not explicitly enforce robustness to domain shifts. The convolutional front-ends used in models like RawNet2 [99] and RawBMamba [104] learn features that are highly discriminative for in-domain data but fail to generalize to novel synthesis artifacts. Even AASIST [106], despite its improved graph attention classifier, is ultimately constrained by its RawNet2-based encoder and shows more failures than those with more specialized architectures.
Wav2Vec-2.0 Feature Extraction Models This category of models demonstrates highly context-dependent performance, achieving State-Of-The-Art results against certain generator families but failing significantly on others. For example, XLSR-AASIST excels on in-domain data like ASVSpoof-21LA ($0.82\%$ EER) but is less effective against modern TTS systems. This sensitivity to generator type is a defining characteristic: performance often collapses against advanced unseen systems (e.g., Matcha-TTS, FireRedTTS-1S) while remaining strong against older GAN or diffusion models. Among these models, XLSR-Conformer+TCM often demonstrates superior score calibration.
@@ 4.3.4 Detector Overall Performance
Conversely, very lightweight models like AASIST ($S_{i}=0.5053$) and RawBMamba ($S_{i}=0.5681$) attained competitive mid-tier scores. Their minimal $\hat{C}_{i}$ penalties compensated for what Section 4.3.3 described as more modest or variable empirical performance (e.g., “limited generalization” or higher $C_{llr}$ for some end-to-end models). Detectors such as RawPCDart ($S_{i}=0.7739$) and RawNet2 ($S_{i}=0.7030$) ranked lower, as their less competitive $P_{i}$ scores were not sufficiently counterbalanced by their model complexities. These $S_{i}$ rankings thus provide a holistic but weight-dependent view, confirming that while the newest generator architectures pose significant detection hurdles (in Section 4.3.2), detectors excelling in broad empirical accuracy can lead when model efficiency is a secondary but still influential consideration.
@@ 5 Future Directions
In this section, we outline concrete and forward-looking research avenues to navigate the escalating race between synthetic voice generation and detection, highlighting needs of both technical innovations and systemic strategies.
@@
6 Related Work
Generalization and Cross Evaluation. Performance often collapses when detectors face unseen generators or corpora [118, 113, 119, 120]. Doan et al. [121] identify highly transferable artifacts from large E2E-GAN generators, and Mishra et al. [122] use Shapley-based explanations to pinpoint shortcut features. Wu et al.’s CodecFake [123] confirms that retraining on neural codec enhances cross-dataset resilience.